Two spherical objects are separated by a distance of 1.63 × 10-3 m. The objects are initially electrically neutral and are very small compared to the distance between them. Each object acquires the same negative charge due to the addition of electrons. As a result, each object experiences an electrostatic force that has a magnitude of 5.54 × 10-21 N. How many electrons did it take to produce the charge on one of the objects
Answer #1

so distance between two objects is 1.63  × 10-3 m

kQ1Q2/r2 =  5.54 × 10-21 N

Q1Q2=1.63 *10-36

Q2 =1.27 *10-18

about 7.56 electrons can carry such charge

so total electrons needed is 15 electrons